OpenStudy (anonymous):

Express the series below in summation notation for the specified number of terms. 7. 4 + 8 + 12 + 16 + 20

OpenStudy (anonymous):

\[\sum_{n=4}^{5}_{?}(n+4)\]

OpenStudy (anonymous):

would this be the equation

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

since u have 5 terms in the series, u need to have the bounds range from 1 to 5, ok ?

ganeshie8 (ganeshie8):

\(\large \sum \limits_{n=1}^5XXX\)

ganeshie8 (ganeshie8):

next look at the series : 4 + 8 + 12 + 16 + 20

ganeshie8 (ganeshie8):

wat kindof series is it ?

OpenStudy (anonymous):

whats the triple x for

ganeshie8 (ganeshie8):

arithmetic or geometric ?

ganeshie8 (ganeshie8):

you need to fill in XXX

OpenStudy (anonymous):

arithmatic

ganeshie8 (ganeshie8):

good, whats the \(n^{th}\) term ?

OpenStudy (anonymous):

n1

ganeshie8 (ganeshie8):

4 + 8 + 12 + 16 + 20 \(n^{th}\) term = \(4 + (n-1)*4 = 4 + 4n - 4 = 4n\)

ganeshie8 (ganeshie8):

plug that in ur sum notation

OpenStudy (anonymous):

so the equation would be \[\sum_{n=1}^{5} 4(n-1)\]

ganeshie8 (ganeshie8):

4 + 8 + 12 + 16 + 20 can be represented in sum notation as : \(\large \sum \limits_{n=1}^5 4n\)